Working on Radicals...
sqrt x^3 = 8: would x be 2? can you please show me step by step? thanks!
3sqrt x^2 = 4: not sure of this one.. the 3 is throwing me; again step by step, it's the only way I can truly understand all this. thanks again!!

Answer by checkley71(8403)

sqrtx~3=8 or squaring both sides we get x~3=64 or x=4
proof sqrt4*4*4=8 sqrt64=8 or 8=8
3sqrtx~2=4 or 3sqrt(x*x)=4 3*x=4 or x=4/3
proof 3sqrt4/3*4/3=4 or 3*4/3=4 or 12/3=4 or 4=4